### v1

# Algorithms for Markov Binomial Chains

Abstract.
We study algorithms to analyze a particular class of Markov population processes that is often used in epidemiology. More specifically, Markov binomial chains are the model that arises from stochastic time-discretizations of classical compartmental models. In this work we formalize this class of Markov population processes and focus on the problem of computing the expected time to termination in a given such model. Our theoretical contributions include proving that Markov binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. We give a PSPACE algorithm for the problem of approximating the time to termination and a direct algorithm for the exact problem in the Blum-Shub-Smale model of computation. Finally, we provide a natural encoding of Markov binomial chains into a common input language for probabilistic model checkers. We implemented the latter encoding and present some initial empirical results showcasing what formal methods can do for practicing epidemilogists.
Key words and phrases:
Markov population processes and formal verification
Work supported by the Flemish inter-university (iBOF) “DESCARTES” project
1. Introduction
We study a class of discrete-time Markov population processes that is often used to model epidemics. In general, Markov population models refers to Markov models whose state space is a discrete partitioning of a population into colonies or compartments. In other words, their set of states $S$ is such that $S\subseteq\mathbb{N}^{k}$ for some $k>0$. Such models arise in the theories of epidemics, population dynamics, rumors, systems biology, and queuing and chemical reaction networks (see, for instance, [Kin69, HJW11, CAdB23] and references therein).
Due to their various important applications, the literature around Markov population models covers interesting mathematical properties of various subclasses. Importantly, and perhaps because of their interest to the systems biology community, there is also a wealth of formal methods to analyze them (see, e.g., [DHSW11, LMW11, BLN18, BBW20, CAdB23]). Importantly, encodings of Markov population models into related stochastic models and language formats accepted by formal-verification tools are known [HJW11].
Most of the literature concerning Markov population processes focuses on continuous-time variants of them. This is in contrast with the class of processes that we study. (Markov) Binomial chains [Bai75] are discrete-time Markov population processes whose transition probabilities are given by a product of probability mass functions of binomial distributions over possible individual transfers between compartments. While this class of model was initially conceived for simple scenarios with small populations, it has recently been used for more complex situations like the analysis of COVID-19 cases in Belgium [AWS+21].
We are unaware of other studies of algorithms and complexity-theoretic results for binomial chains as defined in this work. Related discrete-time stochastic models that have been studied thoroughly include branching processes [ESY17, ESY18, ESY20] and probabilistic vector addition systems [BKKN15, AK23]. Also worth mentioning is an algorithm by Black and Ross to compute the final (population-)distribution for another class of discrete-time Markov population protocols [BR15].
Contributions.
In this work, we initiate the study of algorithms and formal methods for the analysis of binomial chains. We start by formalizing the model in section 2 and then motivating our formalization by giving a self-contained account of how classical compartmental models give rise to binomial chains, in section 3. Then, in section 4, we prove that binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. Finally, in section 5, we give a PSPACE algorithm to approximate the time to termination in a given binomial chain and, in section 6, we also give a direct algorithm for the exact problem (ignoring issues with irrational numbers and the complexity of arithmetic operations). To close the paper, we give an encoding of binomial chains into a common input language for probabilistic model checkers in section 7 and, in section 8, we present some empirical results obtained with an implementation of this encoding.
2. Definitions
We write $\mathbb{N}$ for the set of all natural numbers, including 00; $\mathbb{Q}$ and $\mathbb{Q}_{\geq 0}$, for the sets of all rational and nonnegative rational numbers, respectively; and $\mathbb{R}$ and $\mathbb{R}_{\geq 0}$, for the sets of all real and nonnegative real numbers, respectively. Also, we use $\lg(x)$ to denote the logarithm of $x$ with respect to base $2$; and $\ln(x)$ for the natural logarithm of $x$, i.e. with respect to base $e$. Instead of $e^{x}$, we sometimes write $\exp(x)$.
Let $k\in\mathbb{N}$ be such that $k\geq 1$. We write $[n]$ for the set $\{1,2,\dots,n\}$. For vectors $\boldsymbol{u},\boldsymbol{v}\in\mathbb{Q}^{k}$ with rational entries, we write $\boldsymbol{u}\leq\boldsymbol{v}$ to denote that $u_{i}\leq v_{i}$ for all $1\leq i\leq k$. Let $\ell\in\mathbb{N}$ be such that $\ell\geq 1$ and $\boldsymbol{M}\in\mathbb{Q}^{k\times\ell}$ be a matrix with rational entries. We write $\mathrm{supp}(\boldsymbol{M})$ to denote the support of $\boldsymbol{M}$, that is, the set of indices such that the corresponding entry of $\boldsymbol{M}$ is nonzero. In symbols,
$\mathrm{supp}(\boldsymbol{M})=\{(i,j)\in[k]\times[\ell]:M_{ij}\neq 0\}.$
We extend the notion of support to vectors $\boldsymbol{u}$ in the natural way and write $i\in\mathrm{supp}(\boldsymbol{u})$ instead of $(i,1)\in\mathrm{supp}(\boldsymbol{u})$.
It will also be convenient to have notation for a class of linear transformations from vectors of natural numbers to nonnegative scalars. Let $k\in\mathbb{N}$ be such that $k\geq 1$. We write $\mathbb{L}_{k}$ for the set of all functions $f:\mathbb{N}^{k}\to\mathbb{Q}_{\geq 0}$ such that there exists a vector $\boldsymbol{a}\in\mathbb{Q}^{k}_{\geq 0}$ and a scalar $b\in\mathbb{Q}_{\geq 0}$ for which $f(\boldsymbol{x})=\boldsymbol{a}^{\intercal}\boldsymbol{x}+b$.
Finally, to make sure our notation for them is clear, we give a definition for Markov chains. A discrete-time Markov chain is a tuple $(S,s_{0},P)$ where $S$ is a countable set of states, $s_{0}\in S$ is the initial state, and $P:S\times S\to\mathbb{R}_{\geq 0}$ is a transition probability function, that is, for all $s\in S$ we have that $\sum_{s^{\prime}\in S}P(s,s^{\prime})=1$. We write $X_{t}$, where $t\in\mathbb{N}$, for the random variable representing the state of the chain at the $(t+1)$-th step and $\Pr(X_{t}=s)$ for the probability measure of runs (i.e. sequences of transitions) of the chain with $s$ as their $(t+1)$-th state. In a slight abuse of notation, whenever $S$ is finite, we sometimes write $\boldsymbol{P}$ for the matrix with entries $P_{ij}=P(i,j)$, for all $i,j\in S$.
2.1. Binomial chains
For the rest of this section, let us fix some $k\in\mathbb{N}$ such that $k\geq 1$.
A binomial chain (BC) is essentially a Markov chain $(S,s_{0},P)$ such that $S\subseteq\mathbb{N}^{k}$ and whose transitions correspond to transfers of individuals between compartments modeled by the components of the state vectors. The probability of each individual transfer is based on a binomial distribution whose success probability is a function of the current state.
More formally, a BC $\mathcal{B}$ is a tuple $(\boldsymbol{v},\boldsymbol{T})$ where:
  • $\boldsymbol{v}\in\mathbb{N}^{k}$ is the initial state and
  • $\boldsymbol{T}\in\mathbb{L}_{k}^{k\times k}$ is the transfer matrix.
For intuition, $\boldsymbol{T}$ can be thought of as the adjacency matrix (if we care only about whether a function entry is the zero function or not) of a directed graph with $[k]$ as its vertices. Edges of said graph represent possible transfers of individuals from the source compartment of the edge to the target compartment.
Let us formalize the idea of making $\boldsymbol{T}$ into a (Boolean) adjacency matrix since the notation will be useful later. First, we extend the notion of matrix support to matrices with entries from $\mathbb{L}_{k}$. For $f\in\mathbb{L}_{k}$, we write $f\not\equiv 0$ to denote the fact that $f(\boldsymbol{x})\neq 0$ for some $\boldsymbol{x}\in\mathbb{N}^{k}$. Then, the support of a matrix $\boldsymbol{M}\in\mathbb{L}_{k}^{\ell\times m}$ is defined as follows.
$\mathrm{supp}(\boldsymbol{M})=\{(i,j)\in[\ell]\times[m]:M_{ij}\not\equiv 0\}$
The adjacency matrix is $\left.\boldsymbol{T}\right|_{\mathbb{B}}$ where $(\left.\boldsymbol{T}\right|_{\mathbb{B}})_{ij}=1$ if and only if $(i,j)\in\mathrm{supp}(\boldsymbol{T})$.
The semantics of the BC $\mathcal{B}$ is given via the induced Markov chain $\mathcal{C}_{\mathcal{B}}=(S,s_{0},P)$ where $S=\mathbb{N}^{k}$ and $s_{0}=\boldsymbol{v}$. For the transition probability function $P$, we have that $P(\boldsymbol{u},\boldsymbol{w})>0$ only if there exists $\boldsymbol{M}\in\mathbb{N}^{k\times k}$ such that:
$\mathrm{supp}(\boldsymbol{M})$ ${}\subseteq\mathrm{supp}(\boldsymbol{T})$ (only allowed transfers) (1)
$\forall i\in[k]:\max_{j}M_{ij}$ ${}\leq u_{i}$ (valid binomial outcomes)
and $\boldsymbol{w}$ is the pointwise maximum of $\boldsymbol{0}$ and:
$\boldsymbol{u}+\overbrace{(\boldsymbol{1}^{\intercal}\boldsymbol{M})^{\intercal}}^{\text{incoming transfers}}-\underbrace{\boldsymbol{M}\boldsymbol{1}}_{\text{outgoing transfers}}$
where $\boldsymbol{1}$ is a vector of all ones. In other words, $\boldsymbol{w}$ is such that, for all $j\in[k]$ we have:
$w_{j}=\max\left(0,u_{j}+\sum_{i\in[k]}M_{ij}-\sum_{\ell\in[k]}M_{j\ell}\right).$ (2)
It is easy to see that, for any given $\boldsymbol{u}$, the set of $\boldsymbol{M}$ that satisfy the constraints from Equation 1 is finite. We write $\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ to denote the set of all matrices satisfying Equation 1 and Equation 2. Finally, the transition probability $P(\boldsymbol{u},\boldsymbol{w})$ is defined as follows:
$\sum_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}B(M_{ij};u_{i},1-\exp(-T_{ij}(\boldsymbol{u})))$ (3)
where $B(m;n,p)$ stands for the probability mass function of the binomial distribution:
$B(m;n,p)=\binom{n}{m}p^{m}(1-p)^{n-m}$
with the convention that $10^{0}=$ 1.
Remark 1 (Negative populations).
Note that, in general, the second condition from Equation 1 on matrices $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ is necessary to avoid negative populations but it is not sufficient. This is why $\boldsymbol{a}$ is not always $\boldsymbol{0}$, i.e. why the pointwise maximum with $\boldsymbol{0}$ is needed. A natural stronger condition would be to ask that $\boldsymbol{M}\boldsymbol{1}\leq\boldsymbol{u}$. In order to keep the model as simple as possible and to avoid (further) complicating the transition probability expressions, the simpler (insufficient) condition is often preferred (see, e.g., [AWS+21]). In particular, using the stronger condition would require renormalizing Equation 3 since it would exclude certain outcomes of the binomial distributions.
Having a nonempty set $\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ is not sufficient to guarantee a positive transition probability. This is because the success probability of some binomial distribution may still be 00 for $\boldsymbol{u}$. Excluding that possibility gives us a sufficient and necessary condition.
Lemma 2.
Let $\boldsymbol{u},\boldsymbol{w}\in\mathbb{N}^{k}$. We have $P(\boldsymbol{u},\boldsymbol{w})>0$ if and only if there exists $M\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ such that $\max_{j\in[k]}M_{ij}\leq u_{i}$ for all $i\in[k]$, and $M_{ij}>0$ implies $T_{ij}(\boldsymbol{u})>0$ for all $i,j\in[k]$.
2.2. Interesting subclasses
We now introduce some natural subclasses of binomial chains. The first two are already present in Kingma’s work [Kin69]. Let $(\boldsymbol{v},\boldsymbol{T})$ be a BC.
Simple:
We say that it is simple if for all $i\in[k]$ and all $j\in[k]$, we have that $T_{ij}(\boldsymbol{u})$ can be written as $f(u_{i})$. That is, all transfers from the $i$-th component depend only on the current number of individuals in that compartment.
Closed:
We say that it is closed if all of its transitions preserve the total population. In symbols, for all $\boldsymbol{u},\boldsymbol{w}\in\mathbb{N}^{k}$ we have that:
$P(\boldsymbol{u},\boldsymbol{w})>0\text{ implies }\left\lVert\boldsymbol{u}\right\rVert_{1}=\left\lVert\boldsymbol{w}\right\rVert_{1}.$
Acyclic:
We say it is acyclic if $\left.\boldsymbol{T}\right|_{\mathbb{B}}$ is acyclic. This means that there exists no pair $(i,n)\in[k]\times\mathbb{N}$ such that $(\left.\boldsymbol{T}\right|_{\mathbb{B}})^{n}_{ii}>0$.
Note that being simple and acyclic are properties that can be checked on the description of the BC (concretely, by inspecting the transfer matrix $\boldsymbol{T}$). In contrast, the definition of when a BC is closed seems to depend on the induced Markov chain. Fortunately, we have the following characterization of closed BCs.
Lemma 3.
A BC $(\boldsymbol{v},\boldsymbol{T})$ is closed if and only if its transfer matrix $\boldsymbol{T}$ is such that every row has at most one nonzero entry. That is, |$|\{(i^{\prime},j)\in\mathrm{supp}(\boldsymbol{T}):i=i^{\prime}\}|\leq 1$ for all $i\in[k]$.
Proof 2.1.
If all rows of the transfer matrix have at most one nonzero entry, the stronger condition from 1 coincides with the weaker one from our definition of BC. The contrapositive of the converse is also easy to establish: if the BC is not closed then it must be the case that a vector state with a negative component — say, in dimension $i$ — is reached. In turn, from the definition of the binomial distribution, this can only happen if the $i$-th row of $\boldsymbol{T}$ has at least two nonzero entries.
In the rest of this work, we will mainly study acyclic and closed BCs as they cover interesting models used in epidemiology.
2.3. Computational problems
Suppose $k\in\mathbb{N},k\geq 1$ is fixed and that we are given a BC $\mathcal{B}$ as the tuple $(\boldsymbol{v},\boldsymbol{T})$ with functions from $\mathbb{L}_{k}$ represented as pairs of vectors with rational entries. In turn, assume integers are encoded in binary; and rationals, as pairs of integers representing the numerator and the denominator of the rational number (in reduced form).
The following problems are of practical interest in view of the applications of BCs.
Termination:
asks whether the BC almost surely reaches (final) states $\boldsymbol{t}$ such that, in the induced Markov chain $\mathcal{C}_{\mathcal{B}}$, $P(\boldsymbol{t},\boldsymbol{t})=1$.
Time to termination:
asks to compute the expected number of steps before termination, assuming the BC almost surely terminates.
Regarding the second problem, it is important to note that the value could be irrational. This is because of the exponential function used in the definition of the success probability of the binomial distributions in Equation 3.
In this work, we will primarily focus on algorithms to compute (rational approximations of) the expected time to termination.
3. SIR Models: From ODEs to a bimonial chain
Deterministic SIR Models are simple mathematical models of the spreading of infectious diseases. In them, a population of size $N\in\mathbb{N}$ is partitioned into compartments with labels: $S$ for susceptible, $I$ for infectious, and $R$ for recovered. As illustrated in Figure 1, people may move between compartments following time-dependent dynamics which are usually prescribed by ordinary differential equations (ODEs).
$\dfrac{dS(t)}{dt}$ ${}=-\beta I(t)S(t)$
$\dfrac{dI(t)}{dt}$ ${}=\beta I(t)S(t)-\gamma I(t)$
$\dfrac{dR(t)}{dt}$ ${}=\gamma I(t)$
Hence, we write $S(t)$, $I(t)$, and $R(t)$ to highlight the fact that these values are functions of time $t\in\mathbb{R}_{\geq 0}$. Below, we first state the deterministic SIR model for the case of a closed population [KR08], i.e., births, deaths or infections resulting from contacts with individuals from outside the population are not being considered. (The values $\beta$ and $\gamma$ are explained in the sequel.) Then, we derive a stochastic discrete-time version thereof, based on Bailey’s chain binomial [Bai75].
$S$$I$$R$$\beta$$\gamma$
Figure 1. Overview of the flow of individuals in the SIR model: Following a disease infection, susceptible individuals ($S$) move to an infectious state ($I$) in which they can infect others. Such infectious individuals will recover over time.
3.1. Towards a stochastic SIR model
When an infectious individual makes contact with a susceptible individual, there is some probability that such contact will lead to disease transmission. This probability, multiplied by the contact rate, is denoted by $\beta^{*}$, and we take it to be irrespective of the specific susceptible-infectious pair. Furthermore, we define the force of infection, denoted by $\Lambda(t)$, as the rate for a susceptible individual to become infected at time t. Assuming homogeneous mixing within the population yields the relation
$\Lambda(t)=\dfrac{I(t)}{N}\beta^{*}.$ (4)
To simplify notation, we let $\beta=\beta^{*}/N$. The formulation of the force of infection in (4) is referred to as mass action transmission [KR08]. The exposition above is extended to all susceptible individuals, leading to the continuous-time relation:
$\dfrac{dS(t)}{dt}=-\Lambda(t)S(t)=-\beta I(t)S(t).$ (5)
We start the discretization by fixing $h>0$. By integrating (5) over the time interval $(t,t+h]$, the following recurrence relation is deduced.
[…]
The following integral becomes the cumulative force of infection over $(t,t+h]$.
[…]
By taking the Taylor expansion of (7) around $t$ and considering expressions (4), (6) together, the following holds.
[…]
${}=e^{-\beta(hI(t)+O(h^{2}))}S(t)$
${}\approx e^{-h\beta I(t)}S(t),\text{ for $h$ small}$
The probability $p_{1}(t)=1-e^{-h\beta I(t)}$ will now be regarded as the success probability for the Bernoulli trial corresponding to an interaction between an infectious and a susceptible individual, and the interaction occurs within $(t,t+h]$. By recalling the assumption of homogeneous mixing, the previous discussion suggests we define a random variable $I_{t+h}^{new}$ that follows a binomial distribution and which represents the newly infected individuals.
$I_{t+h}^{new}\sim B(S(t),p_{1}(t)=1-\exp(-h\beta I(t)))$ (8)
To finish this modeling part, the infectious period $\Delta(t)$ is assumed to be exponentially distributed with parameter $\gamma\in\mathbb{R}_{\geq 0}$, irrespective of the individual. Accordingly, the cumulative density function $p_{2}=1-\exp(-h\gamma)$ will represent the probability for an infectious individual to have recovered by time $h>0$. This suggests that the corresponding number of newly recovered individuals $R_{t+h}^{new}$ can be defined as a random variable with binomial distribution as follows.
$R_{t+h}^{new}\sim B(I(t),p_{2}=1-\exp(-h\gamma))$ (9)
3.2. The SIR process
Let $S(0)$, $I(0)$, and $R(0)$ be fixed constants such that $N=S(0)+I(0)+R(0)$ — these are just the initial conditions for the ODE version of an SIR model — and $h>0$. For all $t\in\mathbb{N}$, we define discrete random variables $S_{t}$, $I_{t}$, and $R_{t}$, all of which take values from $\mathbb{N}$. In particular, let $S_{0}=S(0)$, $I_{0}=I(0)$, and $R_{0}=R(0)$. For $t\geq 0$ we base the following definition on Equations (8) and (9):
$S_{t+1}={}$ $S_{t}-Y_{t}$ (10)
$I_{t+1}={}$ $I_{t}+Y_{t}-Z_{t}$ (11)
$R_{t+1}={}$ $R_{t}+Z_{t}$ (12)
with $Y_{t}\sim B(S_{t},1-\exp(-h\beta I_{t}))$ and $Z_{t}\sim B(I_{t},1-\exp(-h\gamma))$.
By definition, we have the property of conservation of population.
Lemma 4.
For all $t\in\mathbb{N}$ we have that $N=S_{t}+I_{t}+R_{t}$.
Let $t\in\mathbb{N}$ be arbitrary and write $(S_{t},I_{t},R_{t})=(m_{1},m_{2},m_{3})$. We will focus on the probability mass function $p^{(t+1)}_{S}$ of $S_{t+1}$. The equations below follow from our definition.
$p^{(t+1)}_{S}(n_{1})={}$ $\Pr(S_{t}=n_{1})$ (13)
[…]
(Here, we again adopt the convention that $10^{0}=$ 1 so that $p^{(t+1)}_{S}(m_{1})$ to be $1$ when $m_{2}=0$.) Similarly, for the probability mass function $p^{(t+1)}_{R}$ of $R_{t+1}$, we get the following equations.
$p^{(t+1)}_{R}(n_{3})={}$ $\Pr(R_{t}=n_{3})$ (14)
[…]
Observe, from (13) and (14), that $S_{t+1}$ and $R_{t+1}$ are independent. Recall that if $X$ and $Y$ are independent then $\Pr(X=x,Y=y)=\Pr(X=x)\Pr(Y=y)=p_{X}(x)p_{Y}(y)$, for $x,y$ in the sample space. Thus, the joint probability mass function $p^{(t+1)}_{S,R}$ of $(S_{t+1},R_{t+1})$ is given by $p^{(t+1)}_{S,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$. Finally, note that $I_{t+1}=N-S_{t+1}-R_{t+1}$, by Lemma 4. Therefore, the joint probability mass function $p^{(t+1)}_{S,I,R}$ of $(S_{t+1},I_{t+1},R_{t+1})$ is also given by $p^{(t+1)}_{S,I,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$.
3.3. A discrete-time Markov chain induced by the binomial chain
Henceforth, we write $X_{t}$ for $(S_{t},I_{t},R_{t})$. For all $t\in\mathbb{N}$, $X_{t}$ is a discrete random variable whose states are vectors from $(n_{1},n_{2},n_{3})\in\mathbb{N}^{3}$ such that $N=n_{1}+n_{2}+n_{3}$ and its probability mass function (as discussed above), is given, for all $t\geq 1$, by:
$p^{(t)}_{S,I,R}(n_{1},n_{2},n_{3})={}$ $\binom{m_{1}}{m_{1}-n_{1}}(\exp(-h\beta m_{2}))^{n_{1}}(1-\exp(-h\beta m_{2})^{m_{1}-n_{1}}$ (15)
${}\binom{m_{2}}{n_{3}-m_{3}}(1-\exp(-h\gamma))^{n_{3}-m_{3}}\exp(-h\gamma)^{m_{2}-n_{3}+m_{3}}$
where $X_{t-1}=(m_{1},m_{2},m_{3})$.
We can now define the transition matrix of the Markov chain. First, note that in Equation (15) the dependency on $t$ can be changed to a dependency on $\boldsymbol{m}=(m_{1},m_{2},m_{3})$. Now, for all states $\boldsymbol{m},\boldsymbol{n}=(n_{1},n_{2},n_{3})$, we define:
$P_{\boldsymbol{m}\boldsymbol{n}}=\begin{cases}p_{S,I,R}^{(\boldsymbol{m})}(n_{1},n_{2},n_{3})&\text{if }n_{1}\leq m_{1}\text{ and }m_{3}\leq n_{3}\leq m_{2}+m_{3}\\ 0&\text{otherwise.}\end{cases}$ (16)
Then, the matrix $\boldsymbol{P}$ is clearly stochastic.
Finally, the transfer matrix $\boldsymbol{T}$ of the BC whose induced Markov chain we defined above, can be inferred from Figure 1 and Equation 15. For concreteness, we give it below.
$\boldsymbol{T}=\begin{pmatrix}0&\boldsymbol{m}\mapsto h\beta m_{2}&0\\ 0&0&\boldsymbol{m}\mapsto h\gamma\\ 0&0&0\end{pmatrix}$
Note that the binomial chain is both closed and acyclic.
To conclude this section we state the following property which follows immediately from the definition of $\boldsymbol{P}$ in Equation 16.
Lemma 5.
Let $\boldsymbol{m}=(m_{1},m_{2},m_{3}),\boldsymbol{n}=(n_{1},n_{2},n_{3})$ be states. Then, $0<\boldsymbol{P}(\boldsymbol{m},\boldsymbol{n})$ if and only if $n_{1}\leq m_{1}$ and $m_{3}\leq n_{3}\leq m_{2}+m_{3}$.
4. Acyclic binomial chains are absorbing
In this section, we recall the notion of absorbing Markov chain. Then, we state interesting properties of such chains that will be useful in the sequel. Finally, we argue that all acyclic BCs induce absorbing Markov chains.
4.1. Absorbing Markov chains
Let $\mathcal{C}=(S,s_{0},P)$ be a Markov chain such that $S$ is finite. We say that state $j\in S$ is reachable from state $i\in S$ if and only if for some $t\in\mathbb{N}$:
$\Pr(X_{t}=j\mid X_{0}=i)>0.$
A state $i\in S$ is absorbing if and only if $P_{ii}=1$. If the set of absorbing states $A\subseteq S$ of the Markov chain $\mathcal{C}$ is not empty and $A$ is reachable from all states, we say $\mathcal{C}$ is an absorbing Markov chain. In an absorbing Markov chain, a state that is not absorbing is transient.
The transition matrix $\boldsymbol{P}$ of an absorbing Markov chain has special properties. Consider an ordering of $S$ such that the $k$ transient states are first, followed by the $\ell$ absorbing states. Now, the transition matrix will have the following canonical form.
$\boldsymbol{P}=\begin{pmatrix}\boldsymbol{Q}&\boldsymbol{R}\\ \boldsymbol{0}&\boldsymbol{I}\end{pmatrix}$ (17)
Above, $\boldsymbol{R}$ is a nonzero $k\times\ell$ matrix; $\boldsymbol{Q}$, a $k\times k$ matrix; and $\boldsymbol{I}$ and $\boldsymbol{0}$, identity and zero matrices, respectively, of the appropriate dimensions.
Proposition 6.
The matrix $\boldsymbol{I}-\boldsymbol{Q}$ has an inverse.
The result above is well known. It is, for example, stated and proven in [GS97, Theorem 11.4]. The inverse of $\boldsymbol{I}-\boldsymbol{Q}$ is commonly written $\boldsymbol{N}$ and called the fundamental matrix. The importance of this matrix will become clear in the next few paragraphs.
4.2. Expected hitting times
Let $A\subseteq S$ be a set of target states. We write $\tau_{A}$ to denote the first hitting time of a state in $A$. Note that $\tau_{A}$ can take countably many values only and they all are nonnegative. Hence, its expectation, denoted $\operatorname{\mathbb{E}}[\tau_{A}]$, satisfies the following (see also [Nor97, Section 1.3]).
$\operatorname{\mathbb{E}}[\tau_{A}]=\sum_{t=0}^{\infty}t\Pr(\tau_{A}=t)+\infty\Pr(\tau_{A}=\infty)$ (18)
Let $i\in S$ be a state and write $k^{A}_{i}$ for the value $\operatorname{\mathbb{E}}[\tau_{A}]$ where the initial state $s_{0}$ of $\mathcal{C}$ is replaced by $i$. The following characterization of the expected hitting times will be useful later. The result is well known and can be found, for instance, in [Nor97, Theorem 1.3.2].
Proposition 7.
The vector of expected hitting times $\boldsymbol{k^{A}}=(k^{A}_{i}:i\in S)$ is the minimal (w.r.t. the product order) nonnegative solution to the following system.
$\begin{cases}k^{A}_{i}=0&\text{if }i\in A\\ k^{A}_{i}=1+\sum_{j\not\in A}P_{ij}k^{A}_{j}&\text{otherwise.}\end{cases}$ (19)
4.3. Expected hitting times in absorbing Markov chains
For absorbing Markov chains, the probability that the process reaches an absorbing state is one. This, in turn, means that the expected hitting times for the set $A\subseteq S$ of absorbing states are always finite. In fact, a formula for the vector of expected hitting times exists in terms of the fundamental matrix (see, e.g. [GS97, Theorem 11.5]).
Proposition 8.
Consider an absorbing Markov chain with absorbing set of states $A\subseteq S$ and fundamental matrix $\boldsymbol{N}$ and let $\boldsymbol{k^{A}}=(k_{i}^{A}:i\in S)$ be the vector of expected hitting times. Then, $(k^{A}_{i}:i\in S\setminus A)=\boldsymbol{N}\boldsymbol{1}$ and $(k^{A}_{i}:i\in A)=\boldsymbol{0}.$
4.4. Acyclic binomial chains are absorbing
We will now argue that every acyclic binomial chain $(\boldsymbol{T},\boldsymbol{v})$ induces an absorbing Markov chain. The summary of our approach is as follows. First, we will define a total order on state vectors. Then, we will establish that the transitions of the induced Markov chain respect this order. Finally, we will also prove that acyclic binomial chains induce finite Markov chains. The claim will follow directly from these properties.
Theorem 9.
Let $\mathcal{B}$ be an acyclic BC. Then, its induced Markov chain $\mathcal{C}_{\mathcal{B}}$ is absorbing.
Since $\left.\boldsymbol{T}\right|_{\mathbb{B}}$ is acyclic, we can use it to sort $[k]$ topologically. That is to say, we can assume that for all $i,j\in[k]$ the following holds.
$T_{ij}\not\equiv 0\implies i<j$ (20)
This means that any $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ will be upper triangular with zeros in the diagonal. Based on this observation, we focus on the lexicographic order on vector states: we write $\boldsymbol{w}\leq_{\mathrm{lex}}\boldsymbol{u}$ if and only if $\boldsymbol{w}=\boldsymbol{u}$ or there exists $j\in[k]$ such that $w_{j}<u_{j}$ and $w_{i}=u_{i}$ for all $1\leq i<j$. We claim that the transitions of the induced Markov chain respect this order.
Lemma 10.
Let $\boldsymbol{u}$ and $\boldsymbol{w}$ be states of the acyclic BC and $P$ the transition probability function of its induced Markov chain. If $0<P(\boldsymbol{u},\boldsymbol{w})$ then $\boldsymbol{w}\leq_{\mathrm{lex}}\boldsymbol{u}$ and $\left\lVert\boldsymbol{w}\right\rVert_{1}\leq\left\lVert\boldsymbol{u}\right\rVert_{1}k$.
Proof 4.1.
First, assume $\boldsymbol{u}\neq\boldsymbol{w}$ as otherwise the claim holds trivially. Let $j^{\prime}\in[k]$ be the index of the first row of $\boldsymbol{M}$ containing some nonzero entry. Hence, we have the following.
$\forall j<j^{\prime}:\sum_{\ell\in[k]}M_{j\ell}=0$
From the triangularity observation above we also get that all columns of $\boldsymbol{M}$ with index $j\leq j^{\prime}$ have only zeros. Hence, we also have the following.
$\forall j\leq j^{\prime}:\sum_{i\in[k]}M_{ij}=0$
It follows from Equation 2 that $w_{j}=u_{j}$ for all $j<j^{\prime}$ and $w_{j^{\prime}}=u_{j^{\prime}}-\sum_{\ell\in[k]}M_{j^{\prime}\ell}$. Since the entries of $\boldsymbol{M}$ are nonnegative and its $j^{\prime}$-th row has some nonzero entry, we get $w_{j^{\prime}}<u_{j^{\prime}}$ and thus $\boldsymbol{w}\leq_{\mathrm{lex}}\boldsymbol{u}$.
For the second property we also rely on $\boldsymbol{M}$ being upper triangular. Because of that fact, we can rewrite Equation 2 as follows for all $j\in[k]$.
$w_{j}$ ${}=\max\left(0,u_{j}+\sum_{i<j}M_{ij}-\sum_{\ell>j}M_{j\ell}\right)$
[…]
${}\leq\sum_{i\leq j}u_{i}$ by Equation 1
It thus follows that $\left\lVert\boldsymbol{w}\right\rVert_{1}\leq\left\lVert\boldsymbol{u}\right\rVert_{1}k$ as claimed.
To conclude the proof of 9 we also argue that acyclic BCs are finite. This follows from 10 and Dickson’s lemma since the lexicographic order is a linear extension of the product order. However, it will be useful later to have an explicit bound on the size of the set of states of the BC. Our bound will rely on 10 and the fact that every state that is reachable in the (Markov chain induced by the) BC is reachable in a small number of steps. As a stepping stone, we prove the claim for the case when the support of intermediate states stays the same.
Lemma 11.
Let $\boldsymbol{u^{(1)}},\dots,\boldsymbol{u^{(n)}}\in\mathbb{N}^{k}$ be vectors with the same support and $t\in\mathbb{N}$ such that $t\geq 1$. If $\Pr(X_{t+n}=\boldsymbol{u^{(n)}},\dots,X_{t}=\boldsymbol{u^{(1)}})>0$ then there exists $\ell<k$ such that $\Pr(X_{t+\ell}=\boldsymbol{u^{(n)}},X_{t}=\boldsymbol{u^{(1)}})>0$.
Proof 4.2.
If $n<k$, the claim holds trivially. Hence, we focus on the case where $n\geq k$.
We will need some additional notation. From the assumptions in the claim, there are matrices $\boldsymbol{M^{(1)}},\dots,\boldsymbol{M^{(n-1)}}$ such that $\boldsymbol{M^{(\ell)}}\in\mathbf{Wit}(\boldsymbol{u^{(\ell)}},\boldsymbol{u^{(\ell+1)}})$ for all $1\leq\ell<n$. In addition, the matrices also satisfy the following for all $1\leq\ell<n$.
$\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}B(M^{(\ell)}_{ij};u^{(\ell)}_{i},1-\exp(-T_{ij}(\boldsymbol{u^{(\ell)}})))>0$
We will now define a second sequence of states $\boldsymbol{w^{(1)}}\dots\boldsymbol{w^{(k)}}$, such that $\boldsymbol{w^{(1)}}=\boldsymbol{u}$ and $\boldsymbol{w^{(k)}}=\boldsymbol{w}$, with their corresponding matrices $\boldsymbol{N^{(1)}},\dots,\boldsymbol{N^{(k-1)}}$. It will be clear that $\boldsymbol{N^{(\ell)}}\in\mathbf{Wit}(\boldsymbol{w^{(\ell)}},\boldsymbol{w^{(\ell+1)}})$, for all $1\leq\ell<k$, based on the triangularity of the $\boldsymbol{M^{(\ell)}}$ and how we will define our $\boldsymbol{N^{(\ell)}}$. Afterwards, we will argue that the corresponding products of the binomial probability mass functions are positive. The intuition is that $\boldsymbol{N^{(\ell)}}$ will correspond to the sum of the $\ell$-th rows of all $\boldsymbol{M^{(\ell^{\prime})}}$ with $\ell^{\prime}>\ell$ and all the rows of $\boldsymbol{M^{(\ell)}}$ strictly below the $\ell$-th one.
We first construct the matrices. For $1\leq\ell<k$ and all $i\in[k]$, we set:
$\boldsymbol{N^{(\ell)}_{i}}=\begin{cases}\boldsymbol{0}&\text{if }i<\ell\\ \sum_{\ell^{\prime}\geq\ell}\boldsymbol{M^{(\ell^{\prime})}_{i}}&\text{if }i=\ell\\ \boldsymbol{M^{(\ell)}_{i}}&\text{otherwise.}\end{cases}$
Now, we define $\boldsymbol{w^{(\ell+1)}}$ as the pointwise maximum of $\boldsymbol{0}$ and $\boldsymbol{w^{(\ell)}}+(\boldsymbol{1}^{\intercal}\boldsymbol{N^{(\ell)}})^{\intercal}-\boldsymbol{N^{(\ell)}}\boldsymbol{1}$.
Let $\ell\in[k]$ be arbitrary. By triangularity of the original witness matrices $\boldsymbol{M^{(\ell^{\prime})}}$ and the definition of the new ones $\boldsymbol{N^{(\ell^{\prime})}}$, we get that $w^{(\ell)}_{i}=w_{i}$ for all $i\leq\ell$ and $w^{(\ell)}_{i}\geq u^{(\ell)}_{i}$ for all $i>\ell$. In particular, this means that $\mathrm{supp}(\boldsymbol{w^{(\ell)}})=\mathrm{supp}(\boldsymbol{u^{(\ell)}})$. Recall that every $T_{ij}$ can be written as $\boldsymbol{x}\mapsto\boldsymbol{a}^{\intercal}\boldsymbol{x}+b$ with $\boldsymbol{a}$ and $b$ being nonnegative. Hence, if $T_{ij}(\boldsymbol{x})>0$ and $\mathrm{supp}(\boldsymbol{y})=\mathrm{supp}(\boldsymbol{x})$, for vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{N}^{k}$, then $T_{ij}(\boldsymbol{y})>0$. We thus conclude that for all $1\leq\ell^{\prime}<\ell$ the following holds.
$\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}B(N^{(\ell^{\prime})}_{ij};w^{(\ell^{\prime})}_{i},1-\exp(-T_{ij}(\boldsymbol{w^{(\ell^{\prime})}})))>0$
Therefore, as required, $\Pr(X_{t+k-1}=\boldsymbol{u^{(n)}},X_{t}=\boldsymbol{u^{(1)}})>0$.
We can now state a concrete bound on the norm of reachable states in an acyclic BC.
Lemma 12.
Let $\mathcal{B}=(\boldsymbol{v},\boldsymbol{T})$ be an acyclic BC. The set $R$ of reachable states in its induced Markov chain is finite. Moreover, for all $\boldsymbol{w}\in R$ we have that $\left\lVert\boldsymbol{w}\right\rVert_{1}\leq\left\lVert\boldsymbol{v}\right\rVert_{1}\exp((\ln k)2^{k^{2}})$.
Proof 4.3.
First, we prove that 11 can be generalized to when the support of intermediate states does change. This will cost us an exponential in terms of $k$. Concretely, we claim that any reachable state $\boldsymbol{w}\in\mathbb{N}^{k}$ can be reached in at most $2^{k^{2}}$ steps. Towards a contradiction, suppose the shortest run witnessing that $\boldsymbol{w}$ is reachable (from $\boldsymbol{v}$) is longer. Then, by pigeonhole principle, the run contains an infix of length at least $k$ such that all states in it have the same support. But then, by 11, we can shorten that infix and obtain a shorter witness.
Second, to obtain the bound on the norm of reachable states $\boldsymbol{w}$ we just need to compose $2^{k^{2}}$ times the bound from 10. We thus obtain:
$\left\lVert\boldsymbol{w}\right\rVert_{1}\leq\left\lVert\boldsymbol{v}\right\rVert_{1}k^{2^{k^{2}}}=\left\lVert\boldsymbol{v}\right\rVert_{1}\exp((\ln k)2^{k^{2}})$
as claimed.
We close this section with the observation that the bound from the previous claim is (asymptotically) good. Indeed, since $k$ is assumed to be fixed, the bound says that the $1$-norm of all reachable states is linear in that of the initial state, i.e. $O(\left\lVert\boldsymbol{v}\right\rVert_{1})$.
5. Approximating the time to termination in polynomial space
In this section, we study the computational complexity of the (approximate) time-to-termination problem for acyclic binomial chains. Somewhat surprisingly, despite the induced Markov chain $\mathcal{C}_{\mathcal{B}}$ of a given BC $\mathcal{B}$ being exponentially larger than the size of the encoding of $\mathcal{B}$ (see 12 and recall the components of $\boldsymbol{v}$ are encoded in binary), we can compute the expected time to termination using polynomial space only. Intuitively, our algorithm consists of a composition of three polynomial-space transducers that have to deal with numbers whose binary encoding may be large (read, exponential). One transducer reads bits of the encoding of the success probabilities from the binomial distributions, another reads bits of the encoding of the transition probabilities, and a last one computes bits of the expected time to termination. However, before we get to that point, we will have to address a small yet important issue: While the input to our computational problems is finite, the transition probabilities of the induced Markov chain may be irrational.
5.1. Rational approximations of induced probabilities
We follow a very simple approach. In short, we use the Taylor expansion of $e^{-x}$ to get a rational approximation thereof. We also derive a bound on the error of our approximation using standard calculus.
Let $x\in\mathbb{Q}_{\geq 0}$. We assume $x$ is given as a pair of positive integers encoding the numerator and denominator, i.e. $x=\frac{a}{b}$ for $a,b\in\mathbb{N}$ with $a,b>0$. We study the function $f(x)=e^{-x}$. The $k$-th order Taylor approximation of $f$ is the following polynomial.
$P_{k}(x)=\sum_{i=0}^{k}\frac{(-x)^{i}}{i!}$
Now, since $-1\leq f^{(k)}(x)\leq 1$ for all $k\in\mathbb{N}$ and all $x\in\mathbb{Q}_{\geq 0}$, the remainder $R_{k}(x)=f(x)-P_{k}(x)$ can be bounded as below using Taylor’s inequality.
$|R_{k}(x)|\leq\frac{x^{k+1}}{(k+1)!}$ (21)
We can now state a sufficient bound on how large $k$ should be to have a small error in our approximation.
Lemma 13.
Let $r\in\mathbb{N}$. If $k\geq 2a^{2}+r$ then $|R_{k}(x)|\leq 2^{-r}$.
Proof 5.1.
We start with a simple observation: There exists $K\in\mathbb{N}$ such that all $k\geq K$ satisfy $x,x^{2}<k$. One can, for instance, take $K=a^{2}$. Now, we claim that for all $k\geq 2K$ the following hold.
$\frac{x^{k}}{k!}<\left(\frac{x}{K}\right)^{k-2K}<1$
The right inequality holds because of our choice of $K$. For the left inequality, we observe the following.
$\frac{x^{k}}{k!}$ ${}<\left(\frac{x}{K}\right)^{k-2K}\frac{x^{2K}}{(2K!)}$ because $\text{because }K^{k-2K}<\frac{k!}{(k-2K)!}$
${}<\left(\frac{x}{K}\right)^{k-2K}\frac{(x^{2})^{K}}{K^{K}}$ $K^{K}<\frac{(2K)!}{K!}<(2K)!$
${}<\left(\frac{x}{K}\right)^{k-2K}$ $\frac{x^{2}}{K}<1$
This means that for $k\geq 2K$ the error decreases exponentially.
It still remains to determine how much larger than $2K$ do we need $k$ to be so that we get an error of at most $2^{-r}$. For this, consider the following.
$\left(\frac{x}{k}\right)^{i}\leq 2^{-r}$ ${}\iff i\lg\frac{x}{K}\leq-r$
${}\iff i\geq\frac{-r}{\lg\frac{x}{K}}$ since $\text{since }x<K$
${}\iff i\geq\frac{r}{\lg K-\lg x}=\frac{r}{\lg a^{2}-\lg\frac{a}{b}}$
${}\iff i\geq\frac{r}{\lg a+\lg b}$
Since the last inequality holds when $i\geq r$, we can choose $k\geq 2a^{2}+r$. The result thus follows by Equation 21.
From the preceding discussion we get a rational approximation of the success probability of the underlying binomial distributions in our model. We still need to determine how to compute it efficiently. It will turn out that we can do so using only polynomial space (in terms of the number of bits required to write $a$, $b$, and $r$ in binary). To be precise, we will be able to query the $i$-th bit of the numerator or denominator (encoded in binary) of the rational approximation.
Lemma 14.
Given $a,b,r,i\in\mathbb{N}$ such that $a,b,r,i>0$, all encoded in binary, we can compute the $i$-th bit of $n$ or $d$ in $\nicefrac{{n}}{{d}}=P_{k}(\nicefrac{{a}}{{b}})$, where $k=2a^{2}+r$, using only space $(\lg(abri))^{O(1)}$.
Before we go into the proof of the claim, some definitions are in order. We write $\mathbf{NC}^{i}$ for the class of decision problems solvable in time $O(\log^{i}n)$, with $n$ the size of the input, on a parallel computer with a polynomial number of processors [Vol99, AB09]. It is known that any problem in $\mathbf{NC}^{i}$ can be solved deterministically using space $O(\lg^{i}(n))$ [Bor77, Theorem 4]. As a concrete example with $i=1$, one can show that computing (a chosen bit of) the sum or product of a list of binary-encoded integers is in $\mathbf{NC}^{1}$ [Vol99, Chapter 1] by exploiting associativity to realize the operations via binary splitting. For an example with $i=2$, we ask, given a list of lists, to compute the sum of the products (of the inner lists). Based on the previous example, this can be done in $\mathbf{NC}^{2}$.
Proof 5.2 (Proof of 14).
Recall the form of the approximation we are proposing.
$P_{k}(\nicefrac{{a}}{{b}})=\frac{n}{d}=\frac{1}{b^{k}k!}\sum_{i=0}^{k}(-a)^{i}b^{k-i}(k-i)!$
This means that $d$ can be taken to be some power of $b$ and the factorial of $k$ while $n$ is a sum of products of powers factorials. From a complexity point of view, the complication is that $k\geq a^{2}$ and $a$ is given in binary, so $n$ and $d$ could require exponentially many bits to represent. (This is also why we focus on determining the value of a single bit only.) Note that computing $n$ and $d$ both amount to computing (sums of) products of at most $k^{3}$ binary-encoded integers. Recall that this problem is in $\mathbf{NC}^{2}$ when $k$ is small, i.e. given in unary. This means it can also be solved using polylogarithmic space, so we conclude, due to $k$ being exponentially large, that we can solve it using polynomial space only.
The argument used above will be repeated two times in the sequel to establish that computing bits of transition probabilities in the induced Markov chain and bits of entries of its fundamental matrix can be done using polynomial space only. Finally, we observe that we can also query bits of an approximation of $1-\exp(-x)$ in polynomial space by following all the same steps while changing $P_{k}(x)$ to remove the first summand and flip all signs.
5.2. Bits of the induced probabilities in polynomial space
Consider a given acyclic binomial chain $(\boldsymbol{v},\boldsymbol{T})$. The main message in this subsection is that we can query bits of the numerator and denominator of Equation 3 (reproduce explicitly below, for convenience) in polynomial space with respect to the encoding size of the given BC. Furthermore, we establish bounds regarding the accumulation of the error because of our usage of a rational approximation of the success probabilities (cf. previous subsection).
$\sum_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}\binom{u_{i}}{M_{ij}}(\overbrace{1-\exp(-T_{ij}(\boldsymbol{u}))}^{\text{computable in {PSPACE}}})^{M_{ij}}(\underbrace{\exp(-T_{ij}(\boldsymbol{u}))}_{\text{this too}})^{u_{i}-M_{ij}}$ (22)
Recall that all integers and rational numbers are encoded in binary. Now, since $\boldsymbol{T}$ is given as an explicit matrix, the product over $(i,j)\in\mathrm{supp}(\boldsymbol{T})$ is small (i.e. a product of a polynomial number of terms w.r.t. the size of the input). The sum over $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ is exponential though, because of the binary encoding. Moreover, the factorials arising from the binomial coefficients are also large since the components of $\boldsymbol{u}$ are given in binary.
We want to argue, as in the previous subsection, that sums of products being in $\mathbf{NC}^{2}$ and thus also in space $O(\lg^{2}(n))$, with $n$ the size of the input, gives us our polynomial-space result. We first need to take care that the numerator and denominator are clearly sums of products or just products. For this, we rewrite Equation 22 as follows, where we are already using the approximation $P_{k}$ from the previous section instead of the (possibly irrational) success probabilities. We write $\tilde{P}$ for this approximate transition probability function.
$\tilde{P}(\boldsymbol{u},\boldsymbol{w})={}$ $\sum_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}\binom{u_{i}}{M_{ij}}\left(\frac{a_{ij}(\boldsymbol{u})}{b_{ij}(\boldsymbol{u})}\right)^{M_{ij}}\left(\frac{c_{ij}(\boldsymbol{u})}{d_{ij}(\boldsymbol{u})}\right)^{u_{i}-M_{ij}}$ (23)
[…]
where $\text{where }\mu\text{ is }\prod_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}b_{ij}(\boldsymbol{u})^{-M_{ij}}d_{ij}(\boldsymbol{u})^{M_{ij}-u_{i}}$ (25)
Above, it is clear that Equation 24 can be used as the numerator of the result and Equation 25 as the reciprocal of the denominator. These are both in the form of a sum of products or a product, as required.
Lemma 15.
Given a BC $(\boldsymbol{v},\boldsymbol{T})$, state vectors $\boldsymbol{u},\boldsymbol{w}\in\mathbb{N}^{k}$, and $r,i\in\mathbb{N}$ (in binary) with $r,i>0$, we can compute the $i$-th bit of $n$ or $d$ in $\nicefrac{{n}}{{d}}=\tilde{P}(\boldsymbol{u},\boldsymbol{w})$ using only polynomial space. Moreover, $|P(\boldsymbol{u},\boldsymbol{w})-\tilde{P}(\boldsymbol{u},\boldsymbol{w})|\leq 2^{-r}$.
Proof 5.3.
The argument to prove $\tilde{P}(\boldsymbol{u},\boldsymbol{w})$ can be computed using polynomial space is the same as the one used in the proof of 14. We just need to make use of Equation 24, Equation 25, and 14 to obtain (exponentially long) sums of products of terms whose bits can be queried in polynomial space. It remains to argue that the error bound holds, and that without having to make $r$ (and thus $k$) in our use of 14 too large.
Consider Equation 23 and note that the only terms with errors are the fractions arising from our approximation of the success probabilities. Multiplying numbers $x,y$ with $10\leq x,y<$ 1 approximated with an error $10<\varepsilon<$ 1 results in at most tripling the error. Similarly, addition results in at most doubling the error (this holds in general though, even if the assumptions stated for multiplication do not hold). Now, by analyzing the exponentiation using repeated squaring, we have at most polynomial tripling of the error in the terms $a_{ij}(\boldsymbol{u})^{M_{ij}}b_{ij}(\boldsymbol{u})^{-M_{ij}}$ and $c_{ij}(\boldsymbol{u})^{u_{i}-M_{ij}}d_{ij}(\boldsymbol{u})^{M_{ij}-u_{i}}$. Now, while the product of these two terms is smaller than $1$, the binomial coefficient is not. In this case, multiplication affects the error much more and we get that it is amplified by at most $\left\lVert\boldsymbol{u}\right\rVert_{\infty}^{\left\lVert\boldsymbol{u}\right\rVert_{\infty}}$. Fortunately, the product with the binomial coefficient once more yields a value smaller than $1$. So, to summarize these observations in symbols, if we started with an error of at most $2^{-r^{\prime}}$ from our use of 13 and 14 then we get that:
$|P(\boldsymbol{u},\boldsymbol{w})-\tilde{P}(\boldsymbol{u},\boldsymbol{w})|\leq\left\lVert\boldsymbol{u}\right\rVert_{\infty}^{\left\lVert\boldsymbol{u}\right\rVert_{\infty}}3^{c}2^{-r^{\prime}}=2^{(\lg\left\lVert\boldsymbol{u}\right\rVert_{\infty})\left\lVert\boldsymbol{u}\right\rVert_{\infty}}3^{c}2^{-r^{\prime}}\leq 2^{\left\lVert\boldsymbol{u}\right\rVert_{\infty}^{2}}3^{c}2^{-r^{\prime}}$
for some small $c\in\mathbb{N}$ — which we could even assume to be encoded in unary. Since $3^{c}4^{-c}\leq 1$, it suffices to use $r^{\prime}=2rc\left\lVert\boldsymbol{u}\right\rVert_{\infty}^{2}$ when appealing to 14 to get the required bounds. Importantly, this means the complexity bound claimed above does hold since $r^{\prime}$ requires only linearly many more bits to be encoded in binary compared to $r$ and $\boldsymbol{u}$.
The last step of our algorithm is arguably the most complex. We intend to compute the fundamental matrix, i.e. the inverse of the induced Markov chain $\mathcal{C}_{\mathcal{B}}$ using polynomial space only. For that, and also as a sanity check, we first wonder how the absolute error $2^{-r}$ can be chosen to be certain it is smaller than the transition probabilities.
Remark 16 (Avoiding small probabilities with large errors).
Let $\boldsymbol{u}$ and $\boldsymbol{w}$ be vector states reachable from $\boldsymbol{v}$ such that $\tilde{P}(\boldsymbol{u},\boldsymbol{w})>0$ and consider the question of determining a sufficient lower bound for $r$ so that $\tilde{P}(\boldsymbol{u},\boldsymbol{w})>2^{-r}$. Ideally $r$ can be encoded in binary using a polynomial number of bits with respect to the rest of the input. To see that this is indeed the case, we can study Equation 22 and find an upper bound for it under the assumption that it is not zero.
$\sum_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}\binom{u_{i}}{M_{ij}}(1-\exp(-T_{ij}(\boldsymbol{u})))^{M_{ij}}(\exp(-T_{ij}(\boldsymbol{u})))^{u_{i}-M_{ij}}$ (26)
[…]
Now, write $n$ for the number of bits used to encode the BC $(\boldsymbol{v},\boldsymbol{T})$. By 12, the number of bits required to encode a state $\boldsymbol{u}$ reachable from $\boldsymbol{v}$ is at most $n+k2^{k^{2}}$. Furthermore, the number of bits required to encode the vectors for the functions $T_{ij}$ is also $n$ so $T_{ij}(\boldsymbol{u})$, for some reachable $\boldsymbol{u}$, satisfies the following when it is nonzero:
$\frac{1}{2^{n+1}}\leq T_{ij}(\boldsymbol{u})\leq 2^{(k+2)n+k2^{k^{2}}}.$
From the above inequalities we get that $\exp(-T_{ij}(\boldsymbol{u}))$ can be made closer to 00 than to $1$ (assuming $T_{ij}(\boldsymbol{u})$ is not 00). Together with Equation 26 we get that if $\tilde{P}(\boldsymbol{u},\boldsymbol{w})$ is not zero then it satisfies the following.
$\tilde{P}(\boldsymbol{u},\boldsymbol{w})>\exp(-2^{(k+2)n+k2^{k^{2}}})^{k^{2}}\geq 2^{-2k^{2}2^{(k+2)n+k2^{k^{2}}}}$ (27)
Hence, by choosing $r$ larger than $2k^{2}2^{(k+2)n+k2^{k^{2}}}$, which can be encoded in a polynomial number of bits in $n$ (because $k$ is fixed), we get the desired inequality.
5.3. Time to termination in polynomial space
Given, a nonsingular $n\times n$ matrix $\boldsymbol{A}$ with $n$-bit integer entries (encoded in binary), we must output its inverse $\boldsymbol{A}^{-1}$ in the form of a pair $(\mathrm{adj}(\boldsymbol{A}),\det(\boldsymbol{A}))$ consisting of the adjugate and the determinant of $\boldsymbol{A}$. It is known that computing chosen bits of the numerator or denominator of a chosen entry of $A^{-1}_{ij}=\nicefrac{{\mathrm{adj}(\boldsymbol{A})_{ij}}}{{\det(\boldsymbol{A})}}$ is in $\mathbf{NC}^{2}$ [Coo85, Proposition 5.2].
Recall that the set of reachable states in the Markov chain induced by a given BC $(\boldsymbol{v},\boldsymbol{T})$ is finite, yet exponential (due to the binary encoding of the integers) 12. Now, we would like to appeal to 8 in combination with the above observations to obtain a polynomial-space algorithm for obtaining the bits of the numerator and denominator of a chosen $k^{A}_{i}$ from the induced Markov chain.
Theorem 17.
Given a BC $(\boldsymbol{T},\boldsymbol{v})$ and $r,i\in\mathbb{N}$ in binary with $r,i>0$, we can compute the $i$-th bit of $n$ or $d$ in $\nicefrac{{n}}{{d}}=\tilde{k}^{A}_{\boldsymbol{v}}$ using only polynomial space. Moreover, $|\tilde{k}^{A}_{\boldsymbol{v}}-k^{A}_{\boldsymbol{v}}|\leq 2^{-r}$.
Proof 5.4.
We first reorder the states to get $\boldsymbol{P}$ in its canonical form (see Equation 17). We rely on the lexicographic ordering introduced in subsection 4.4 to achieve this. Concretely, if we want the $i$-th state $\leq_{\mathrm{lex}}$-smaller than $\boldsymbol{v}$, we can enumerate all vector states satisfying the bound from 12 to find the maximal state $\boldsymbol{w}$ smaller than $\boldsymbol{v}$ and repeat $i-1$ times from $\boldsymbol{w}$, all using polynomial space only. Further note that checking whether a state $\boldsymbol{w}$ is absorbing can be implemented in polynomial space using 15 by enumerating all states not equal to $\boldsymbol{w}$ and confirming the probability to transition to them is 00. (This last check can even be done without having to approximate the success probabilities!)
Before applying the complexity result for inverting a matrix, we need to deal with the fact that we want to invert a matrix of rational numbers and not one of integers. The natural approach would be to factor out a common denominator (as we have done in previous subsections). However, because of the shape of $\boldsymbol{I}-\boldsymbol{Q}$ in our case, this is not necessary. By our choice of reordering of the states and 10, we have that $\boldsymbol{I}-\boldsymbol{Q}$ is upper triangular. In turn, this means that its inverse $\boldsymbol{N}=\nicefrac{{\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})}}{{\det(\boldsymbol{I}-\boldsymbol{Q})}}$ and thus also $\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})$ are upper triangular. Since the nonzero entries of the the adjugate $\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})$ are obtained as signed determinants of minors of $\boldsymbol{I}-\boldsymbol{Q}$ obtained by removing rows $\boldsymbol{u}$ and columns $\boldsymbol{w}$ with $\leq_{\mathrm{lex}}\boldsymbol{w}$. Such minors will necessarily be upper triangular too. Therefore, their determinants are just products of entries of $\boldsymbol{I}-\boldsymbol{Q}$. Importantly, no sum is needed, so we can focus independently on numerators or denominators of the entries of $\boldsymbol{I}-\boldsymbol{Q}$ based on whether we want bits of the numerator or denominator of an entry of $\boldsymbol{N}$.
From the discussion above and the $\mathbf{NC}^{2}$ bound for matrix inversion (and the polynomial-space algorithms from previous subsections) that the bits of fundamental matrix $\boldsymbol{N}$ can be queried using polynomial space, we still need to establish the same for $\boldsymbol{N}\boldsymbol{1}$. This is not a problem since multiplying by $\boldsymbol{1}$ on the right amounts to adding rows of $\boldsymbol{N}$, which can be done in $\mathbf{NC}^{1}$ if $\boldsymbol{N}$ is given explicitly. Altogether, inverting a matrix and then adding its rows can be done in $\mathbf{NC}^{3}$ thus also sequentially while using only $O(\lg^{3}(n))$ space, with $n$ the size of the input. This means that in our (exponentially large) induced Markov chain, the bits of the (numerator and denominator) of $\boldsymbol{N}\boldsymbol{1}$ can be queried in polynomial space. When using the approximation of the success probabilities, we write $\boldsymbol{\tilde{k}^{A}}$ instead of $\boldsymbol{k^{A}}$.
We have already argued why the complexity bound in 17 holds. It remains for us to prove that the error bounds are true. First, observe that we can ensure the error of the entries of $\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})$ and $\det(\boldsymbol{I}-\boldsymbol{Q})$ is small. We do this by choosing the error $\varepsilon=2^{-r^{\prime}}$ allowed in our application of 15 exponentially smaller than the required one, i.e. $r$, to compensate for the increase in the error due to the multiplication of entries from $\boldsymbol{I}-\boldsymbol{Q}$ to obtain each entry. This can be done without affecting our complexity analysis since we can encode $r^{\prime}$ using at most polynomially more bits than what was used to encode $r$. (See also the error-propagation analysis in the proof of 12 where we did almost the same). Now, since the right multiplication by $\boldsymbol{1}$ to get $\boldsymbol{N}\boldsymbol{1}$ can also be dealt with similarly, the only remaining complication is the error accumulated by division in computing $\nicefrac{{\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})}}{{\det(\boldsymbol{I}-\boldsymbol{Q})}}$. Write $x$ and $y$ for the approximated numerator and denominator of any entry of the matrix $\nicefrac{{\mathrm{adj}(\boldsymbol{I}-\boldsymbol{Q})}}{{\det(\boldsymbol{I}-\boldsymbol{Q})}}$ and $a$ and $b$ for the actual values. Assuming we choose $r^{\prime}$ as indicated in 16 so that $0\leq\nicefrac{{\varepsilon}}{{a}},\nicefrac{{\varepsilon}}{{b}}\leq 1$, the following hold:
$\frac{x}{y}={}$ $\frac{a\pm\varepsilon}{b\pm\varepsilon}=\frac{a}{b}\left(\frac{1\pm\frac{\varepsilon}{a}}{1\pm\frac{\varepsilon}{b}}\right)$
[…]
and therefore, $|\nicefrac{{x}}{{y}}-\nicefrac{{a}}{{b}}|$ is the absolute value of the last expression above minus $\nicefrac{{a}}{{b}}$. We thus get the following inequalities.
$\left|\frac{x}{y}-\frac{a}{b}\right|={}$ $\left|\frac{\varepsilon}{b}+\left(\frac{a\pm\varepsilon}{b\pm\varepsilon}\right)\left(\frac{\varepsilon}{b}\right)\right|$
[…]
We can now establish a lower bound for $b$ much like that of Equation 27. In symbols:
$b\geq 2^{-2^{n^{c}}}$
holds, for some constant $c\in\mathbb{N}$. The last inequality above can now be rewritten as follows.
$\left|\frac{x}{y}-\frac{a}{b}\right|\leq{}$ $\frac{\varepsilon}{b}\left(\frac{2}{b-\varepsilon}\right)\leq\frac{2^{-r^{\prime}}}{2^{-2^{n^{c}}}}\left(\frac{2}{2^{-2^{n^{c}}}-2^{-r^{\prime}}}\right)$
[…]
This means that if $r^{\prime}\geq r+4(2^{n^{c}})$ then our approximation has an absolute error of at most $2^{-r}$, exactly as required. Since this can be achieved using polynomially more bits than that needed to encode $r$, our complexity result holds too.
6. Time to termination for SIR models in the Blum-Shub-Smale model
This section is meant as a more pragmatic approach to computing the expected time to termination. For this, we focus on the concrete SIR binomial chain. Furthermore, we change from our familiar Turing-machine model of computation to a Blum-Shub-Smale (BSS) machine [BSS88]. In a BSS machine, we have registers that can hold arbitrary real numbers (including irrational ones) and applying rational operations on them takes a single time step. Finally, we also assume that $\exp(-h\beta)$ and $\exp(-h\gamma)$ are given as part of the input. We will show that, in this context, the expected time to termination can be computed in time polynomial with respect to the $1$-norm of the initial vector.
Let us write $Q$ for the set of states of the Markov chain induced by the given SIR binomial chain $(\boldsymbol{v},\boldsymbol{T})$, so $Q\subseteq\{0,1,\dots,N\}^{3}$, where $N=\left\lVert\boldsymbol{v}\right\rVert_{1}$, and the components of a state $\boldsymbol{m}\in Q$ correspond to susceptible, infectious, and recovered, in that order. We also write $A\subseteq Q$ for its absorbing states, thus $A=\{\boldsymbol{m}\in Q\mid m_{2}=0\}$ by 5. Note that, since SIR binomial chains are closed, one of the components of the vector states is redundant and $|Q|\leq N^{2}$. Hence, the dimensions of $\boldsymbol{P}$ (the matrix representation of the transition probability function of the induced Markov chain) are at most $N^{2}$.
Remark 18 (A first algorithm, polynomial in N).
Recall that the vector $\boldsymbol{k^{A}}$ of expected hitting times can be computed using the expression in 7. This already gives us a naive algorithm that runs in time polynomial in $N$ if $\boldsymbol{P}$ is given. Indeed, Gaussian elimination requires only a cubic number of arithmetic operations. Therefore, given $\boldsymbol{P}$, we can compute $\boldsymbol{N}=(\boldsymbol{I}-\boldsymbol{Q})$ in time $O(N^{6})$.
6.1. A second algorithm, now without Gauss
In light of the proof of 17, one may wonder whether Gaussian elimination is needed to obtain $\boldsymbol{N}$. As we will briefly show, it can in fact be avoided in favor of back substitution because here too $\boldsymbol{P}$ and thus $\boldsymbol{I}-\boldsymbol{P}$ are upper triangular.
Theorem 19.
Given an SIR BC $(\boldsymbol{T},\boldsymbol{v})$, we can compute $\boldsymbol{k^{A}}$ in time $O(N^{4})$ in the Blum-Shub-Smale model.
Before proving the result, we observe that 5 gives us a characterization of the nonzero terms in $\boldsymbol{P}$. In this way, for SIR binomial chains, the expression from 7 is reduced to what is shown in the next lemma.
Lemma 20.
For all states $\boldsymbol{m}\in Q$, we have that:
$k^{A}_{\boldsymbol{m}}=\begin{cases}0&\text{if }m_{2}=0,\\ 1+\sum_{n_{1}=0}^{m_{1}}\sum_{n_{3}=m_{3}}^{m_{2}+m_{3}}P(\boldsymbol{m},\boldsymbol{n})k^{A}_{\boldsymbol{n}}&\text{otherwise,}\end{cases}$
where $\boldsymbol{n}=(n_{1},N-n_{1}-n_{3},n_{3})$.
Our present goal is to order the (vector) states according to a total order as we did in subsection 4.3. The result above suggests a possible order to achieve this. In words, we will use the colexicographic ordering. Since SIR binomial chains are closed, the strict version of the order is defined as follows.
$(m_{1},m_{2},m_{3})<_{\mathrm{colex}}(n_{1},n_{2},n_{3})\iff m_{3}<n_{3}\text{or }m_{3}=n_{3}\text{ and }n_{1}<m_{1}.$
We can now sort $Q$ based on $<_{\mathrm{colex}}$ so that $\boldsymbol{P}$ is in its canonical form (see Equation 17). By 8, we have the following relation.
$(\boldsymbol{I}-\boldsymbol{Q})\boldsymbol{k^{A}}=\boldsymbol{1}$ (28)
Hence, due to our choice of order on the states, together with 5, we get that $\boldsymbol{I}-\boldsymbol{Q}$ is upper triangular. This enables the computation of $\boldsymbol{k^{A}}$ by means of back substitution, which requires only a quadratic number of arithmetic operations.
Lemma 21.
Given an SIR BC $(\boldsymbol{v},\boldsymbol{T})$, we can compute $\boldsymbol{k^{A}}$ in time $O(N^{4})$ in the Blum-Shub-Smale model, assuming $\boldsymbol{P}$ is given.
It remains to argue that $\boldsymbol{P}$ can be precomputed in $O(N^{4})$. Unfortunately, computing every entry of $\boldsymbol{P}$ using Equation 15 seems to require time $O(N)$. Indeed, while exponentiation can be realized using a logarithmic number of operations via iterated squaring, we are not aware of an algorithm to compute factorials using a sublinear number of multiplications. [^1] We do know that, in the Turing-machine model of computation, one can obtain seemingly better bounds [Bor85]. However, those depend on the complexity of multiplying numbers being a function of the amount of bits used to encode them.
6.2. Precomputation of the transition probabilities using dynamic programming
Recall Equation (15) gives us that for all states $\boldsymbol{m}\neq\boldsymbol{n}$ such that $n_{1}\leq m_{1}$ and $m_{3}\leq n_{3}\leq m_{2}+m_{3}$ the probability $P(\boldsymbol{m},\boldsymbol{n})$ of transitioning from $\boldsymbol{m}$ to $\boldsymbol{n}$ is as follows.
$\binom{m_{1}}{m_{1}-n_{1}}\exp(-h\beta m_{2})^{n_{1}}(1-\exp(-h\beta m_{2}))^{m_{1}-n_{1}}$ (29)
$\binom{m_{2}}{n_{3}-m_{3}}(1-\exp(-h\gamma))^{n_{3}-m_{3}}\exp(-h\gamma)^{m_{2}-n_{3}+m_{3}}$ (30)
The crux of our algorithm is the following recurrence.
Lemma 22.
Let $\boldsymbol{m},\boldsymbol{n}$ be states such that $n_{1}<m_{1}$ and $m_{3}<n_{3}\leq m_{2}+m_{3}$. Then, $P(\boldsymbol{m},\boldsymbol{n})=\alpha(\boldsymbol{m},\boldsymbol{n})P(m_{1}-1,m_{2},m_{3}+1,\boldsymbol{n}),$ where:
$\alpha(\boldsymbol{m},\boldsymbol{n})=\frac{m_{1}(m_{2}-n_{3}+m_{3}+1)(1-\exp(-h\gamma))(1-\exp(-h\beta m_{2}))}{(n_{3}-m_{3})(m_{1}-n_{1})\exp(-h\gamma)}.$
Proof 6.1.
Note that $P(\boldsymbol{m},\boldsymbol{n}),P(m_{1}-1,m_{2},m_{3}+1,\boldsymbol{n})>0$ by 5 and our assumptions. We first consider the binomial coefficients of (29) and (30).
$\binom{m_{1}}{m_{1}-n_{1}}\binom{m_{2}}{n_{3}-m_{3}}$
[…]
Importantly, since we have assumed that $m_{3}<n_{3}$ and $n_{1}<m_{1}$, the denominator of the fraction above is not 00. On the other hand, for the terms involving “probabilities” — that is, exponential terms — we observe that the following is true for (29).
$\exp(-h\beta m_{2})^{n_{1}}(1-\exp(-h\beta m_{2}))^{m_{1}-n_{1}}$
[…]
Similarly, for the probabilities in (30) we note the following.
$(1-\exp(-h\gamma))^{n_{3}-m_{3}}\exp(-h\gamma)^{m_{2}-n_{3}+m_{3}}$
[…]
Indeed, $\alpha(\boldsymbol{m},\boldsymbol{n})$ is exactly the product of the coefficients on the right-hand sides of the above equalities. This concludes the proof as the product of the left-hand sides is exactly $P(\boldsymbol{m},\boldsymbol{n})$ and that of the right-hand sides (after factoring out $\alpha$) is $P(m_{1}-1,m_{2},m_{3}+1,\boldsymbol{n}$).
Based on 22, we propose the following algorithm.
{algo}
An efficient algorithm to compute $P(\boldsymbol{m},\boldsymbol{n})$ for all $\boldsymbol{m}$ and $\boldsymbol{n}$
1:for $m_{1}=0,\dots,N$ do
2:     $X(\boldsymbol{m},\boldsymbol{m})=1$, with $\boldsymbol{m}=(m_{1},0,N-m_{1})$
3:end for
4:for $M=1,\dots,N$ do
5:     for $m_{2}=1,\dots,M$ do
6:         $\boldsymbol{m}\leftarrow(M-m_{2},m_{2},N-M)$
7:         for $n_{3}=m_{3},\dots,m_{2}+m_{3}$ do $\triangleright$ Fix $n_{1}=m_{1}$
8:              $X(\boldsymbol{m},\boldsymbol{n})\leftarrow P(\boldsymbol{m},\boldsymbol{n})$, with $\boldsymbol{n}=(m_{1},N-m_{1}-n_{3},n_{3})$
9:         end for
10:         for $n_{1}=0,\dots,m_{1}$ do $\triangleright$ Fix $n_{3}=m_{3}$
11:              $X(\boldsymbol{m},\boldsymbol{n})\leftarrow P(\boldsymbol{m},\boldsymbol{n})$, with $\boldsymbol{n}=(n_{1},N-n_{1}-m_{3},m_{3})$
12:         end for
13:         for $n_{3}=m_{3}+1,\dots,m_{2}+m_{3}$ do
14:              for $n_{1}=0,\dots,m_{1}-1$ do
15:                  $\boldsymbol{n}\leftarrow(n_{1},N-n_{1}-n_{3},n_{3})$
16:                  $\boldsymbol{m^{\prime}}\leftarrow(m_{1}-1,m_{2},m_{3}+1)$
17:                  $X(\boldsymbol{m},\boldsymbol{n})\leftarrow\alpha(\boldsymbol{m},\boldsymbol{n})X(\boldsymbol{m^{\prime}},\boldsymbol{n})$
18:              end for
19:         end for
20:     end for
21:end for
We observe that 6.1 clearly terminates because all for-loops are bounded and there are no jump statements in the code. Regarding its time complexity, we note that we only ever nest at most $4$ for-loops and $P(\boldsymbol{m},\boldsymbol{n})$ is never used (i.e. computed naively in $O(N)$ steps) within a for-loop nesting of depth $4$.
Proposition 23 (Complexity).
The worst-case time complexity of 6.1 is $O(N^{4})$ in the Blum-Shub-Smale model.
For correctness, we have the following statement.
Proposition 24 (Correctess).
Let $X$ be as computed by 6.1. Then, $X(\boldsymbol{m},\boldsymbol{n})=P(\boldsymbol{m},\boldsymbol{n})$ for all pairs of states $\boldsymbol{m},\boldsymbol{n}$ such that $n_{1}\leq m_{1}$ and $m_{3}\leq n_{3}\leq m_{2}+m_{3}$.
Proof 6.2.
First, one needs to take care that the previous transition probabilities must be defined before they are being used. Note that this is only relevant in lines 1319. It thus follows from the order in which the states $\boldsymbol{m}$ are traversed that $X(\boldsymbol{m^{\prime}},\boldsymbol{n})$ has already been computed in a previous iteration.
Now, it is easy to see that $X(\boldsymbol{m},\boldsymbol{n})=P(\boldsymbol{m},\boldsymbol{n})$ if either $m_{1}=n_{1}$ or $m_{3}=n_{3}$ since lines 8 and 11 literally assign the right-hand side to the left-hand side of the equality. If neither equality holds, then $X(\boldsymbol{m},\boldsymbol{n})$ is computed in the loop 1319 and $X(\boldsymbol{m},\boldsymbol{n})=P(\boldsymbol{m},\boldsymbol{n})$ by 22. Indeed, the conditions that $n_{1}<m_{1}$ and $m_{3}<n_{3}$ are guaranteed by values over which $n_{3}$ and $n_{1}$ range in the loops from lines 13 and 14.
19 follows from the results above, together with 21.
7. Time to termination using probabilistic model checkers
In this section, we take a step back from SIR binomial chains and move to the general class of acyclic BCs. It seems difficult to generalize our algorithms for SIR binomial chains, especially since we exploited the property of them being closed. Instead, we propose to encode the computation of the time to termination as an instance of a probabilistic model checking problem. In particular, we aim at a small encoding (i.e. without explicitly constructing the induced Markov chain) that does not depend on the exact initial vector state. This is desirable to avoid reencoding the original BC every time the initial population changes. We target the PRISM input language [KNP11] for our encoding since it is supported by state-of-the-art tools such as PRISM itself, Storm [HJK+22], and Modest [HH14].
Let us fix an acyclic BC $(\boldsymbol{v},\boldsymbol{T})$. As in the previous section, we will avoid the problem of irrational probabilities by asking that they be part of the input. More precisely, we will assume that for all linear functions $T_{ij}=\sum_{\ell=1}^{k}a_{\ell}x_{\ell}+a_{0}$ from $\boldsymbol{T}$ we have precomputed the values $p_{ij\ell}=e^{-a_{\ell}}$. These are then stored using some finite precision and taken to be part of the input. Using the new notation, we can rewrite Equation 3 as follows.
$\sum_{\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})}\prod_{(i,j)\in\mathrm{supp}(\boldsymbol{T})}\binom{u_{i}}{M_{ij}}\left(1-p_{ij0}\prod_{\ell=1}^{k}p_{ij\ell}^{u_{\ell}}\right)^{M_{ij}}\left(p_{ij0}\prod_{\ell=1}^{k}p_{ij\ell}^{u_{\ell}}\right)^{u_{i}-M_{ij}}$ (31)
In the next section, instead of focusing on the exact syntax of the PRISM language, we introduce a new model we call stochastic counter machines. Encoding such machines into a PRISM program is straightforward for someone familiar with the language. Additionally, this intermediate reduction will allow us to highlight the difficulties of attempting a direct encoding of binomial chains into PRISM.
7.1. Stochastic counter machines
A stochastic counter machine (or SCM) in dimension $k$ is a tuple $\mathcal{M}=(Q,q_{0},\boldsymbol{c_{0}},T)$ where $Q$ is a finite set of (control) states, $q_{0}\in Q$ is the initial state, $\boldsymbol{c_{0}}\in\mathbb{N}^{k}$ is the initial value of the counters (to which we sometimes refer using Greek letters), and $T$ is a finite set of tuples $(q,g,u,q^{\prime},p)$ where: $q,q^{\prime}\in Q$, the guard $g$ is a matrix-vector pair $\boldsymbol{A}\in\mathbb{Q}^{n\times d},\boldsymbol{b}\in\mathbb{Q}^{n}$, the udapte $u$ is also a matrix-vector pair $\boldsymbol{U}\in\mathbb{Q}^{d\times d},\boldsymbol{r}\in\mathbb{Q}^{d}$, and $p\in\mathbb{Q}_{\geq 0}$.
A configuration of the SCM is a pair $(q,\boldsymbol{c})\in Q\times\mathbb{N}^{k}$ — we often write $q(\boldsymbol{c})$ instead of $(q,\boldsymbol{c})$. The transition $(q,g,u,q^{\prime},p)$, with $g=(\boldsymbol{A},\boldsymbol{c})$ and $u=(\boldsymbol{U},\boldsymbol{r})$, is enabled from a configuration $q(\boldsymbol{c})$ if and only if $\boldsymbol{A}\boldsymbol{c}\leq\boldsymbol{b}$. Further, the $t$-successor of $q(\boldsymbol{c})$ is the configuration $q^{\prime}(\boldsymbol{U}\boldsymbol{c}+\boldsymbol{r})$ with probability $p$.
We need some constraints on SCMs to ensure configuration-successor pairs are unique and that a proper distribution is defined. For all configurations $q(\boldsymbol{c})$, we ask that:
  • For all transitions $t_{1},t_{2}$, if both $t_{1}$ and $t_{2}$ are enabled from $q(\boldsymbol{c})$ and the $t_{1}$- and $t_{2}$-successors of $q(\boldsymbol{c})$ are the same then $t_{1}=t_{2}$.
  • Let $t_{1},t_{2},\dots$, where $t_{i}=(q,g_{i},u_{i},q^{\prime}_{i},p_{i})$, be an enumeration of the transitions enabled from $q(\boldsymbol{c})$. Then, we have that $\sum_{i\geq 1}p_{i}=1$.
The semantics of the SCM $\mathcal{M}$ is given, like that of a BC, via an induced Markov chain $\mathcal{C}_{\mathcal{M}}=(S,s_{0},P)$ where $S=Q\times\mathbb{N}^{k}$ and $s_{0}=q_{0}(\boldsymbol{c_{0}})$. For the transition probability function $P$, we have that $P(q(\boldsymbol{c}),q^{\prime}(\boldsymbol{c^{\prime}}))=p>0$ if and only if $q^{\prime}(\boldsymbol{c^{\prime}})$ is the $t$-successor of $q(\boldsymbol{c})$, for some transition $t$, with probability $p$. The constraints above ensure that $P$ is well defined.
$q_{0}$$q_{1}$$q_{2}$$[\chi_{3}>0]$$1-p:\begin{pmatrix}[l]\chi_{1}&\leftarrow\chi_{1}\\ \chi_{2}&\leftarrow\chi_{2}\\ \chi_{3}&\leftarrow\chi_{3}-1\end{pmatrix}$$p:\begin{pmatrix}[l]\chi_{1}&\leftarrow\chi_{1}\\ \chi_{2}&\leftarrow\chi_{2}+1\\ \chi_{3}&\leftarrow\chi_{3}-1\end{pmatrix}$$1:\begin{pmatrix}[l]\chi_{1}\leftarrow\chi_{1}\\ \chi_{2}\leftarrow 0\\ \chi_{3}\leftarrow\chi_{1}\end{pmatrix}$$[\chi_{3}=0]$
Figure 2. A stochastic counter machine that simulates a binomial distribution. Control states are depicted as circles and transitions $(q,g,u,q^{\prime},p)$ as arrows from $q$ to $q^{\prime}$. Guards are shown in square brackets along a transition, together with probability-update pairs $p:(u)$ with $p$ being the probability and $u$ the update. If the guard is trivial (i.e. it is satisfied by all configurations) then we omit it; if the update is the identity, we also omit it; and if the probability is $1$, again we omit it. Finally, we group transitions from a common state and with the same guard using a solidly filled circle with the guard shown before the circle and the individual probabilities and updates shown after.
7.2. Binomial and Bernoulli stochastic counter machines
The SCM depicted in Figure 2, encodes a binomial distribution with success probability $p$. Intuitively, from the initial configuration $q_{0}(c_{1},c_{2},c_{3})$ the machine simulates $c_{1}$ Bernoulli trials, each with success probability $p$. The third counter $\chi_{3}$ is used to count from $c_{1}$ (copied when transitioning from $q_{0}$ to $q_{1}$) to 00 by decrementing on both transitions from $q_{1}$ to itself. The right transition from $q_{1}$ to itself simulates a success and thus increments the value of the second counter, where we store the total number of successful Bernoulli trials. The left transition leaves the value of the second counter untouched since it corresponds to a failed Bernoulli trial.
From the discussion above, it should be clear that the SCM implements a binomial distribution in the sense that the distribution over successor configurations with control state $q_{2}$ is given exactly by the probability mass function of that distribution.
Lemma 25.
Let $\mathcal{M}$ be the SCM from Figure 2 and consider its induced Markov chain $\mathcal{C}_{\mathcal{M}}$. For all $t,t^{\prime},c_{1},c_{2},c_{3}\in\mathbb{N}$ with $t^{\prime}-t\geq c_{1}$ it holds that:
$\Pr(X_{t^{\prime}}=q_{2}(c_{1},m,0)\mid X_{t}=q_{0}(c_{1},c_{2},c_{3}))=B(m;c_{1},p).$
The reader may have already started realizing our intention: We want to use binomial SCMs to simulate the innnermost products from Equation 31 which correspond to binomial distributions. Unfortunately, unless $p_{ij\ell}$ is $1$ for all $\ell>0$, the SCM we just presented is not good enough because it simulates Bernoulli trials with a constant success probability $p$. To resolve this, we present a second SCM that allows us to model a Bernoulli trial where the success probability is a function of the counter values of the initial configuration. More precisely, the function is $1-p_{0}\prod_{\ell=1}^{k}p_{\ell}^{c_{k}}$, where $p_{\ell}\in\mathbb{Q}_{\geq 0}$ and $p_{\ell}\leq 1$ for all $0\leq\ell\leq k$.
The SCM depicted in Figure 3, encodes a parametric Bernoulli distribution. From the initial configuration $q_{0}(c_{1},c_{2},c_{3},c_{4})$, and much like in the binomial SCM, the last counter is used as a temporary counter while the second-to-last counter holds the result of the trial: $c_{3}$ if it fails and $c_{3}+1$ if it succeeds. For instance, on the transition from $q_{0}$ to $q_{1}$, the last counter copies the value $c_{1}$. Then, while its value is nonzero, there is a transition back to $q_{1}$ with probability $p_{1}$ that decrements the counter. If its value has reached zero, the machine transitions to $q_{2}$ while setting the value of the last counter to $c_{2}$. There, it again loops on the same state and decrements the counter while its value is nonzero, this time with probability $p_{2}$. On both loops, the alternative transitions reach state $q_{x}$. Clearly, the probability of eventually reaching the state $q_{y}$ is $p_{0}\prod_{\ell=1}^{2}p_{\ell}^{c_{\ell}}$. Hence, the probability of eventually reaching $q_{x}$ is $1-p_{0}\prod_{\ell=1}^{2}p_{\ell}^{c_{\ell}}$. Now, for $\ell>2$, the SCM can be extended further below $q_{2}$ by adding copies of the same structure mutatis mutandis.
$q_{0}$$q_{1}$$q_{2}$$q_{x}$$q_{y}$$p_{0}:\begin{pmatrix}[l]\chi_{4}\leftarrow\chi_{1}\end{pmatrix}$$[\chi_{4}>0]$$[\chi_{4}>0]$$p_{1}:\begin{pmatrix}[l]\chi_{4}\leftarrow\chi_{4}-1\end{pmatrix}$$1-p_{1}:\begin{pmatrix}[l]\chi_{3}\leftarrow\chi_{3}+1\\ \chi_{4}\leftarrow 0\end{pmatrix}$$[\chi_{4}=0](\chi_{4}\leftarrow\chi_{2})$$1-p_{2}:\begin{pmatrix}[l]\chi_{3}\leftarrow\chi_{3}+1\\ \chi_{4}\leftarrow 0\end{pmatrix}$$[\chi_{4}=0]$$p_{2}:\begin{pmatrix}[l]\chi_{4}\leftarrow\chi_{4}-1\end{pmatrix}$
Figure 3. A stochastic counter machine that simulates a parametric Bernoulli distribution with $\ell=2$. Most notation is as in Figure 2, but we also omit trivial updates.
Lemma 26.
Let $\mathcal{M}$ be an SCM as in Figure 3 for $\ell\in\mathbb{N}$ and consider its induced Markov chain $\mathcal{C}_{\mathcal{M}}$. For all $\boldsymbol{c}\in\mathbb{N}^{k+2}$ and all $t,t^{\prime}\in\mathbb{N}$ with $t^{\prime}-t\geq\left\lVert\boldsymbol{c}\right\rVert_{1}$ it hold that:
$1-\Pr(X_{t^{\prime}}=q_{x}(c_{1},\dots,c_{k},c_{k+1}+1,0)\mid X_{t}=q_{0}(\boldsymbol{c}))$
[…]
$q_{0}$$q_{1}$$q_{3}$$q_{2}$$q_{4}$$q_{5}$$1:\begin{pmatrix}\alpha_{0}\leftarrow s\end{pmatrix}$$[\alpha_{0}>0]:\begin{pmatrix}[l]\alpha_{1}\leftarrow i\\ \alpha_{0}\leftarrow\alpha_{0}-1\end{pmatrix}$$[\alpha_{0}=0]$$[\alpha_{1}>0]$$[\alpha_{1}=0]$$1-p_{0}:\begin{pmatrix}\chi_{0}\leftarrow\chi_{0}+1\end{pmatrix}$$p_{0}:\begin{pmatrix}\alpha_{1}\leftarrow\alpha_{1}-1\end{pmatrix}$$1:\begin{pmatrix}\alpha_{1}\leftarrow i\end{pmatrix}$$[\alpha_{1}=0]$$1:\begin{pmatrix}[l]s\leftarrow s-\chi_{0}\\ i\leftarrow i+\chi_{0}-\chi_{1}\\ r\leftarrow r+\chi_{1}\\ \chi_{0}\leftarrow 0\\ \chi_{1}\leftarrow 0\\ \end{pmatrix}$$[\alpha_{1}>0]$$p:\begin{pmatrix}\chi_{1}\leftarrow\chi_{1}+1\\ \alpha_{1}\leftarrow\alpha_{1}-1\end{pmatrix}$$1-p_{1}:\begin{pmatrix}\alpha_{1}\leftarrow\alpha_{1}-1\end{pmatrix}$
Figure 4. The SCM we construct for the SIR binomial chain from section 3
7.3. The final construction and going beyond reachability
We can suitably compose parametric Bernoulli SCMs with and binomial SCMs to simulate a BC. After a series of binomial SCMs (to simulate the innermost product from Equation 31), we have at most $k^{2}$ counters whose values correspond to the number of individuals that are to be transferred for this simulated transition of the BC. As an example, we give the full construction for SIR binomial chains in Figure 4. The SCM uses $7$ counters, to which we will refer as $s,i,r,\alpha_{0},\alpha_{1},\chi_{0},\chi_{1}$. Let $p_{0}=1-\exp(-h\beta)$ and $p_{1}=1-\exp(-h\gamma)$ and observe that $\beta_{0}\sim B(s,p_{0}^{i})$ and $\beta_{1}\sim B(i,p_{1})$. as expected. Hence, going once from $q_{0}$ to itself simulates exactly one timestep in the binomial chain.
So far, the formal claims establishing correctness of our construction speak only about transition and eventual reachability probabilities. We can also encode the expected time to termination by adding rewards to the transitions of the SCM. To be precise, we can formalize rewards as pairs $(g,r)$ of guards and rational rewards. Then, all transitions taken from a configuration $q(\boldsymbol{c})$ that satisfies the guard yield a reward of $r$ [^2] To the expert reader: Yes, this is exactly a reward structure in the PRISM language. For our purpose, the reward function should be $+1$ for each simulated step of the binomial chain before reaching an absorbing state, e.g. on the transition from $q_{0}$ to $q_{1}$, and 00 for all other transitions. To make sure only transitions before reaching the absorbing states get a reward, we can add one counter that keeps track of the sum of all counter values corresponding to components of the BC which influence the value of some $T_{ij}$. Then, this sum will be 00 if and only if the current (simulated) vector state is absorbing.
The next claim summarizes the properties of the constructed SCM. There, we subindex probability functions to highlight the probability space of the (induced) Markov chain to which they belong. Furthermore, we refer to the total number of states and counters of an SCM as its size.
Theorem 27.
For all acyclic BC $\mathcal{B}$ we can construct an SCM $\mathcal{M}$ with rewards, of size $k^{O(1)}$, and an injection $\mu$ from states of $\mathcal{B}$ to configurations of $\mathcal{M}$ so that the following hold.
  • There exists $m\in\mathbb{N}$ such that for all $t,t^{\prime},n\in\mathbb{N}$, with $t<t^{\prime}$ and $m<n$, and all states $\boldsymbol{u},\boldsymbol{w}$ of $\mathcal{B}$ we have $\Pr_{\mathcal{B}}(X_{t^{\prime}}=\boldsymbol{w}\mid X_{t}=\boldsymbol{w})=\Pr_{\mathcal{M}}(X_{t^{\prime}+n}=\mu(\boldsymbol{w})\mid X_{t}=\mu(\boldsymbol{u})).$
  • Moreover, the expected time to reach $\boldsymbol{w}$ from $\boldsymbol{u}$ in $\mathcal{C}_{\mathcal{B}}$ is the expected sum of rewards before reaching $\mu(\boldsymbol{w})$ from $\mu(\boldsymbol{u})$ in $\mathcal{C}_{\mathcal{M}}$.
8. Experimental Results
We implemented our translation from binomial chains to stochastic counter machines in a prototype tool we called Inform. More concretely, Inform translates files in an explicit representation of the transfer matrix of a binomial chain into into a PRISM-language file. Using it and a probabilistic model checker such as Storm [HJK+22], we can verify properties of the encoded binomial chain. The second model is a simplified COVID-19 model based on [AWS+21]. In this section we do precisely this. Below, we introduce the binomial chain we studied and the properties we checked.
[…]
Figure 5. Overview of the flow of individuals in the (single-age group version of the) binomial chain for the early stages of COVID-19 in Belgium from [AWS+21]
8.1. The Belgian COVID-19 binomial chain
[…]
All of the above results in the linear function $T_{SE}$, i.e. the entry of the transfer matrix $\boldsymbol{T}$ corresponding to the indices for the compartments $S$ and $S$, being nontrivial. In fact, all other entries of $\boldsymbol{T}$ are constants while $T_{SE}$ depends on around 400400~{}400400 components of the current vector state. (We refer the interested reader to [AWS+21] for the ODEs and the binomial-chain formulation of the time-discretized compartmental model and to its supplementary materials for the values of all constants.) For our experiments, we simplified the COVID-19 binomial chain to summarize all age groups back into a single compartment.
Note that the binomial chain, is acyclic, but not closed. This seems more like a quirk of the modelling formalism rather than a desirable feature in the context of the time bounds studied by the authors of [AWS+21] which span part of the COVID-19 epidemic and some months after it. This, paired with the observation that binomial chains seem to allow the possibility of spontaneous infection of all susceptible individuals in a single transition, inspired our choice of properties to check for in the model.
PopInc:
How likely is it that the population does not remain constant?
OS:
How likely is it that a given portion of the population moves from one compartment to another, in one shot?
EoE:
What is the expected time before the end of the epidemic?
For the OS property, we chose to focus on when this happens along the very first transition of the binomial chain. We give our exact encoding of these properties as given to the model checkers in Table 1.
8.2. Experimental setup
[…]
All experiments were run on a cluster where each node had an Intel(R) Xeon(R) Platinum 8168 CPU @ 2.70GHz with 64GiB of memory and no GPU. We ran Storm from the movesrwth/storm:stable docker container and, based on a number of local experiments, chose the sparse engine for all numbers reported henceforth. Storm’s version was 1.7.1. For Modest, we used version v3.1.237-g2f62162c7. We denote time-outs with TO: we stopped the computation after 1 hour (for small populations); memory-outs with MO: the program was terminated because it ran out of memory. All code and scripts to re-run the experiments can be found here: https://github.com/UA-FOTS/inform.
Table 1. The properties of interest stated in PRISM-style probabilistic computation tree logic and in terms of the Belgian COVID-19 model. The PopInc property states: What is the probability (P=?) that we avoid an error state (that is only reached when the population is not preserved) until the infectious compartments are depleted. When these are empty, it is impossible for more people to become infected. Note that for the OS property, we on the start of the simulation: What is the probability that we avoid infecting susceptible individuals until we infect all of them?
Property Probabilistic Temporal Logic Formula
[…]
EoE R{time_step}=? [F $(E+I_{presym}+\dots+I_{icu}=0)$]
8.3. Results
For both the population increase (PopInc) as well as the end of epidemic (EoE) property, we see Storm running out of memory already for populations of 10. The one-shot (OS) property performs really well in Storm. This may be because our formalization of the OS property only checks the first time individuals change compartment. Storm seems to be taking this into account when building the state-space. Modest performs significantly better than Storm for the PopInc property. Moreover, the size of the population has a much smaller impact on the run-time compared to Storm. However, for the EoE property, we see that Modest struggles. In order to not time out for small instances, the width of the confidence interval of Modest was set to $0.9$ and, even in this case, runtimes were significantly higher than for the PopInc property.
Table 2. Performance of Storm compared to Modest on the COVID-19 model. For Storm, we used default parameters and the sparse engine. For the EoE property, Modest was run with max run length 0 and width 0.9; for the others, with max run length 0 and width 0.01.
Property Population Storm Modest
[…]
PopInc (2, 1, 1, 1) 2.916 3.4s
(3, 1, 1, 1) 21.201 6.9s
(4, 1, 1, 1) 297.23s 5.9s
(5, 1, 1, 1) 2352.066s 5.7s
(6, 1, 1, 1) 14756.769s 4.5s
(7, 1, 1, 1) MO 4.3s
EoE (2, 1, 1, 1) 3.400s 1023.5s
(3, 1, 1, 1) 27.314s 997.8s
(4, 1, 1, 1) 570.862s 1069.3s
(5, 1, 1, 1) 5325.083s 1048.3s
(6, 1, 1, 1) 42751.95 1039.0s
(7, 1, 1, 1) MO 1080.8s
OS (2, 1, 1, 1) 0.123s 0.2s
(3, 1, 1, 1) 0.125s 0.1s
(4, 1, 1, 1) 0.141s 0.2s
(5, 1, 1, 1) 0.186s 0.1s
(6, 1, 1, 1) 0.195s 0.1s
(7, 1, 1, 1) 0.223s 0.1s
(8, 1, 1, 1) 0.244s 0.2s
(9, 1, 1, 1) 0.274s 0.1s
(10, 1, 1, 1) 0.306s 0.2s
Finally, and based on the results summarized above, we used Modest to analyze the COVID-19 model with realistic populations. The results are shown in Figure 6. We observe that the run-time of Modest grows almost linearly with respect to $S$.
[…]
It is worth mentioning that we did compare the translation-based solution against our custom-optimized algorithm from section 6 for SIR binomial chains and the EoE property. Even for small populations, our algorithm from section 6 outperformed both model checkers.
9. Conclusion and future work
We have started the study of binomial chains through the lens of formal methods. In this work, we provided two main theoretical results. First, we established that acyclic binomial chains almost surely terminate. Second, we proved the the problem of approximating the time to termination is in PSPACE and gave a direct algorithm for the exact problem (ignoring the complexity of arithmetic and the annoyances of irrational numbers). We also provided a more pragmatic approach in the form of an encoding into the PRISM language. For this last approach we also presented some experiments. Based on the empirical results, we can conclude that state-of-the-art probabilistic model checkers are not (yet) powerful enough to deal with epidemiological models like the one proposed in [AWS+21].Indeed, while Modest is capable of handling simple probabilistic queries for realistic populations, it still seems to struggle with quantitative queries such as the expected end of epidemic property. In this direction, more research is needed to find, for instance, good abstractions.
Finally, there are natural decision problems for binomial chains that we did not consider in this work. For instance, the following, based on values studied in [AWS+21], seem relevant.
Finite-horizon peak:
asks to compute the maximal expected population in a given compartment within a finite horizon (given in binary).
Finite-horizon accumulated population:
asks to compute the total expected population having been in a given compartment within a finite horizon (given in binary).